OpenStudy (anonymous):

HELP!!! Use the substitution method to solve the following system of equations. 3x + 5y = 3 x + 2y = 0

OpenStudy (shamil98):

x = -2y so. 3(-2y) + 5y = 3 -6y + 5y = 3 -y = 3 y = -3 x = -2(-3) x = 6 (6,-3) is your solution,

OpenStudy (anonymous):

In the below system, solve for y in the first equation. x + 3y = 6 2x - y = 10